OpenStudy (anonymous):

The probability that a battery will last 10 hr or more is 0.8, and the probability that it will last 15 hr or more is 0.11. Given that a battery has lasted 10 hr, find the probability that it will last 15 hr or more

OpenStudy (anonymous):

What do you know about conditional probability and Bayes Theorem?

OpenStudy (anonymous):

not much. Does that have anything to do with the problem?

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

put A = the battery lasts 10 hours or more, so \(P(A)=.8\) and B = the batter lasts 15 hours or more so \(P(B)=.11\)

OpenStudy (anonymous):

by definition \(P(A|B)=\frac{P(A\cap B )}{P(B)}\) so we only need \(P(A\cap B)\)

OpenStudy (anonymous):

but since \(B\subset A\) we know \(P(A\cap B)=P(B)\) so your answer is \(\frac{P(B)}{P(A)}\)

OpenStudy (anonymous):

sorry i messed up with the A and B you want \[P(B|A)=\frac{P(A\cap B)}{P(A)}\]

OpenStudy (anonymous):

to make a long story short your answer is \[\frac{.11}{.8}\]

OpenStudy (callisto):

What I've been thinking is that probability = ''desired outcome'' / ''possible outcome'' In this case, desired outcome have a probability of 0.11 possible outcome is the given condition, that is 0.8 Put them back, then I'll get P required = 0.11/0.8. Actually, I don't understand the third step satellite73 has written. My probability sucks.

OpenStudy (anonymous):

satallite73 was right. Thank you!